OpenStudy (anonymous):

The probability of Samantha scoring an A on a geology test (event A) is 0.46, and the probability of Judith scoring an A on the same test (event B) is 0.54. The probability of both Samantha and Judith scoring an A is 0.2484. Given this information, which statement is true?

OpenStudy (anonymous):

which statement?

OpenStudy (anonymous):

since \(.54\times .46=.2484\) the events are independent, so i guess they are not cheating off each other do you have statements to check?

OpenStudy (ybarrap):

Don't know your statements but the probability that at least one of them gets an A is $$ \large{ P(A~OR~B)= P(A)+P(B)-P(A~AND~B)\\ =0.46+0.54-0.2484=0.7516 } $$